Calculate the wavelength of the electromagnetic radiation required to excite an electron from the ground state to the level with n = 4 in a one-dimensional box 37.7 pm in length.
Answer #1

The formula for the energy of the particle in a 1D box is given by below equation

E=n2h2/8ma2

Where n= energy levels

h= planks constant

m=mass of the particle

a=length of the box

in the given problem

n=4, for this we have to calculate the energy associated with n=6 and ground state that is n=1

then \DeltaE=E4-E1

\DeltaE=(n42h2 / 8ma2) –(n12h2 / 8ma2)

      = h2(n42-n12) / 8ma2

As we know that

E=hc/\lambda ( where h = planks constant; c= velocity of light and \lambda = wavelength

So wavelength = hc/E

                           = hc / h2(n42-n12) / 8ma2

                             = 8ma2c / h (n42-n12)

                             = (8) (9.109 X 10-31 Kg) (0.449 X 10-10 m)2 (3 X 10-8 m/s) / (6.626 X 10-34 J s) (16-1)

                            = 8 X 9.109 X 0.449 X 0.449 X 3 X 10-43 / 6.626 X 15 X 10-34

                            = 44.07 X 10-43 / 99.39 X 10-34

                             = 0.4434 X 10-9 m

                               = 434.4 pm

answer= 434.4 pm